Use the following equation to complete parts (a) through (c) below.

          Markup (margin) percent

  1. Find the markup percent given the following.

     Retail = $85.00; Cost = $30.00

  1. Find the retail price given the following.

          Cost = $35.00; Markup (margin) percent = 60%

  1. Find the cost given the following.

          Retail = $99.00; Markup (margin) percent = 50%
Answer #1

WE KNOW THAT
MARKUP = NEW PRICE - INITIAL PRICE
MARKUP % = MARKUP VALUE / INITIAL PRICE * 100
MARKDOWN = INITIAL PRICE - NEW PRICE
MARKDOWN % = MARKDOWN VALUE / INITIAL PRICE * 100


1. MARKUP
85 - 30 = 55
MARKUP % = 55 / 30 * 100 = 183.33

2. RETAIL PRICE = INITIAL PRICE + MARKUP PERCENTAGE
= 35 + (.60 * 35) = 56

3. NOW MARKUP PERCENTAGE IS MORE THAN THE INITIAL VALUE AND THEREFORE, 50% MARKUP VALUE WOULD BE 150 PERCENT THE INITIAL VALUE:

COST = FINAL PRICE / PERCENTAGE INCREASE = 99/1.50 = $66